# Which Way Does Time Flow? A Psychophysics-Grounded Evaluation for Vision–Language Models

### v2

@@ 6.1 Motion Category Analysis
Category-specific patterns. The largest F1 deficits appear in Proceed and Fall, where all models struggle to capture the causal and gravitational cues defining temporal direction. Diffusion and Put show moderate F1 scores but remain below human performance even in forward playback, suggesting incomplete modeling of dispersion and goal-directed manipulation. Division is comparatively easier for both humans and models, likely due to clear temporal discontinuities (e.g., splitting). Notably, Cosmos-Reason1-7B achieves relatively strong backward F1, particularly on Fall and Diffusion, consistent with its explicit AoT-oriented training, whereas Gemini-2.5-Pro and GPT-4.1 perform better in forward playback but drop sharply when the temporal order is reversed. Overall, these results suggest that current VLMs rely primarily on directional visual priors and dataset correlations rather than robust physical causality understanding.

### v3

@@ 5.3 Reasoning Effort Ablation
5.4 SFT Performance
We fine-tuned Qwen2-VL-7B and Qwen2.5-VL-7B on 1,000, 3,000, and 5,000 training examples to test whether explicit supervision improves AoT performance (Table 11). Although fine-tuning mitigated the extreme forward bias seen in vanilla models (in the 3000 example setting, Qwen2-VL-7B Backward F1: 0→48.2, Qwen2.5-VL-7B Backward F1: 19.5→43.4), it did not lead to real task competence: accuracy stayed near chance (~50%) across all training set sizes. Performance plateaued after 1,000 examples, with no gains from additional data. These results suggest that AoT task might require different training approaches beyond conventional SFT.
5.5 FPS Ablation
@@ 6.1 Motion Category Analysis
Category-specific patterns. The largest F1 deficits appear in Proceed and Fall, where all models struggle to capture the causal and gravitational cues defining temporal direction. Diffusion and Put show moderate F1 scores but remain below human performance even in forward playback, suggesting incomplete modeling of dispersion and goal-directed manipulation. Division is comparatively easier for both humans and models, likely due to clear temporal discontinuities (e.g., splitting). Notably, Cosmos-Reason1-7B achieves relatively strong backward F1, particularly on Fall and Division, consistent with its explicit AoT-oriented training, whereas Gemini-2.5-Pro and GPT-4.1 perform better in forward playback but drop sharply when the temporal order is reversed. Overall, these results suggest that current VLMs rely primarily on directional visual priors and dataset correlations rather than robust physical causality understanding.
@@ 8 Limitations
Appendix A Appendix
A.1 Instructions in Manual CoT Experiments
The instruction used in the Simple CoT setting is shown in Figure 6. The instruction for the Multi-step CoT setting is shown in Figure 7.
(Instruction) Is this video played forward (F) or backward (B)? Pay attention to moving parts in the video as they may be the visual clues to indicate the direction of the playback. Think step by step before giving out your final answer. Finish your answer with F for forward or B for backward only.
Figure 6: The instruction for the Simple CoT setting.
Is this video played forward or backward?
Think in three steps: Observation, Assumption, and Conclusion.
In Observation, describe what you see in the video honestly, without making any early assumptions about whether it is played forward or backward. Pay close attention to moving elements such as peoples movements, the flow of liquid, or any changes in the size or position of objects. In Assumption, make a reasonable guess about what should normally happen in this scene if it were played forward. Base this guess on common sense, including causal relationships, the laws of physics, and the arrow of time. In Conclusion, check whether your Observation matches or contradicts your Assumption. If what you observed follows the expected behavior based on your assumption, then the video is played forward. If what you observed goes against your assumption, then the video is being played backward.
Finish your answer with F for forward or B for backward only.
Figure 7: The instruction used for the Multi-step CoT setting.